OpenStudy (anonymous):

Select the counterexample that makes this conjecture false: For any real number x, if x2 ≥1, then x ≥1.

hartnn (hartnn):

take any negative number less than 1 say x=-2 x^2=(-2)^2=4 so x^2=4 > 1 but x= -2 < 1

OpenStudy (anonymous):

so the answer comes to x=2

hartnn (hartnn):

no,here u just need to give an counterexample..... so i gave an example when x=-2

hartnn (hartnn):

are there options?

OpenStudy (anonymous):

yes x=1,x=1/4,x=2,x=-3

hartnn (hartnn):

so there's x=-3 x^2=(-3)^2=9 >1 x=-3 <1

OpenStudy (anonymous):

so it would be x=1

OpenStudy (anonymous):

can you gave the exact sol.l. of this problem

hartnn (hartnn):

how 1 ? i showed u a counterexample with x=-3.

hartnn (hartnn):

its x=-3

OpenStudy (anonymous):

by differntial equatin solution